Campbell Supper Co. paid a $0.662 dividend per share in 2013, which grew to $0.80 in 2016. This growth is expected to continue.

What is the value of this stock at the beginning of 2017 when the required return is 8.3 percent? (Round the growth rate, g, to 4 decimal places. Round your final answer to 2 decimal places.)
Answer #1

The growth rate of dividends is :

The dividend paid in 2013 = $0.662 and the dividend paid in 2016 is $0.8. So, the growth rate of dividends is ,

0.662* (1 + g)^3 = $0.8

So, the growth rate is :

=( 0.8/ 0.662)^0.33

= 6.5143%

So, the value of the stock in 2017 is :

P 0 = D1/ Re -g

= 0.8 * (1.0651) / 0.0179

= $47.60 ( rounded off to two decimal places)